OpenStudy (anonymous):

Can someone please explain to me about vectors? I really dont understand and want to be able to do it myself. medal and fan if helped.Thanks!

OpenStudy (anonymous):

Mathematics. a quantity possessing both magnitude and direction, represented by an arrow the direction of which indicates the direction of the quantity and the length of which is proportional to the magnitude. Compare scalar (def 4). such a quantity with the additional requirement that such quantities obey the parallelogram law of addition. such a quantity with the additional requirement that such quantities are to transform in a particular way under changes of the coordinate system. any generalization of the above quantities. 2. the direction or course followed by an airplane, missile, or the like. 3. Biology. an insect or other organism that transmits a pathogenic fungus, virus, bacterium, etc. any agent that acts as a carrier or transporter, as a virus or plasmid that conveys a genetically engineered DNA segment into a host cell. 4. Computers. an array of data ordered such that individual items can be located with a single index or subscript. verb (used with object) 5. Aeronautics. to guide (an aircraft) in flight by issuing appropriate headings. 6. Aerospace. to change direction of (the thrust of a jet or rocket engine) in order to steer the craft.

OpenStudy (sleepyjess):

What specifically about vectors?

OpenStudy (anonymous):

how to solve for them . i know nothing of them

OpenStudy (anonymous):

for example one of the questions is : A spotlight is mounted on a wall 7.4 feet above the floor in an office building.it is used to light a door 9.3 feet from the wall.To the nearest degree what is the angle of depresssion from the spotlight to the bottom of the door?

OpenStudy (anonymous):

i looked it up and the answer was 39 but i need to know how to solve for it

OpenStudy (anonymous):

A vector is a mathematical concept that has both magnitude and direction. Detailed explanation of vectors may be found at the Wikibooks module Linear Algebra/Vectors in Space. In physics, vectors are used to describe things happening in space by giving a series of quantities which relate to the problem's coordinate system. A vector is often expressed as a series of numbers. For example, in the two-dimensional space of real numbers, the notation (1, 1) represents a vector that is pointed 45 degrees from the x-axis towards the y-axis with a magnitude of \sqrt 2 . Commonly in physics, we use position vectors to describe where something is in the space we are considering, or how its position is changing at that moment in time. Position vectors are written as summations of scalars multiplied by unit vectors. For example: a \hat{i} + b \hat{j} + c \hat{k} where a, b and c are scalars and \hat{i}, \hat{j} and \hat{k} are unit vectors of the Cartesian (René Descartes) coordinate system. A unit vector is a special vector which has magnitude 1 and points along one of the coordinate frame's axes. This is better illustrated by a diagram. A vector itself is typically indicated by either an arrow: \vec{v}, or just by boldface type: v, so the vector above as a complete equation would be denoted as: \vec{v} = a \hat{i} + b \hat{j} + c \hat{k} The magnitude of a vector is computed by |\vec{v}| = \sqrt{\sum_i(x_i^2)}. For example, in two-dimensional space, this equation reduces to: |\vec{v}| = \sqrt{x^2+y^2}. For three-dimensional space, this equation becomes: |\vec{v}| = \sqrt{x^2+y^2+z^2}.